OpenStudy (anonymous):

Find the equation of the line tangent to the surface at the given point P and in the direction of the given unit vector U.

OpenStudy (anonymous):

\[f(x,y)=x ^{2}-y ^{2} at P=(2,3), u=<\frac{ \sqrt{2} }{ 2 },\frac{ \sqrt{2} }{ 2 }>\]

OpenStudy (anonymous):

I know I have to take the partial derivative, which I got 2x and -2y

OpenStudy (anonymous):

I also know that I have to insert P into F(x,y). I got -5 with that